OpenStudy (anonymous):

John solved x^2 = 6x as follows: x^2/x = 6/x x=6 Did he get all the possible solutions to the equation? Explain the steps in solving the equation correctly.

OpenStudy (shamim):

bt it has 2 solution because its a quadrutic equation

OpenStudy (shamim):

equation has a term as x^2

OpenStudy (anonymous):

as x^2 = 6x x^2 - 6x = 0 x(x-6) = 0 x = 0, x = 6 are the possible solutions

OpenStudy (anonymous):

thank you so much i was just thinking the same thing! :)